OpenStudy (anonymous):

4푣 − 13 − 10푣 = 12 − 24푣 + 5

OpenStudy (inkyvoyd):

Hey, I can't read that - my browser is unicode. Could you please replace the "푣" with the corresponding chars?

OpenStudy (anonymous):

ohh okei, its 4v − 13 − 10v = 12 − 24v + 5

OpenStudy (inkyvoyd):

Alrght, the first step is to isolate the variable. How would we do that?

OpenStudy (anonymous):

so it's -6v - 13 = 17 - 24v

OpenStudy (inkyvoyd):

Yes. Now we get v on one side

OpenStudy (anonymous):

-30v -13 = 17 ?

OpenStudy (inkyvoyd):

Careful there. -6v-13+24v=17-0v

OpenStudy (anonymous):

oh...

OpenStudy (inkyvoyd):

anyways, keep going :)

OpenStudy (anonymous):

and then 18v - 13 = 17 - 0v

OpenStudy (inkyvoyd):

Yes. Now isolate the variable - move the 13 to the other side

OpenStudy (anonymous):

18v - 30 = -0v

OpenStudy (inkyvoyd):

Uhh, you'll want to do this actually 18v=17-0+13

OpenStudy (anonymous):

oh oki.

OpenStudy (inkyvoyd):

btw, -0v is just 0, because zero times anything is 0

OpenStudy (anonymous):

that makes 18v = 30

OpenStudy (inkyvoyd):

Yes. Now divide both sides by 18 to get rid of the 18 on the left side.

OpenStudy (inkyvoyd):

Then, write the fraction in lowest terms.

OpenStudy (anonymous):

5 / 3

OpenStudy (inkyvoyd):

yup!